If 150cm

3

150cm3 of steam at 100

C

100∘C is added to 250cm

3

250cm3 water at 40

C

40∘C, how much steam(in cm

3

cm3) is condensed into water at 100

C

100∘C?


1
Expert's answer
2020-12-06T17:17:00-0500
msR=mwc(TT0)ρsVsR=ρwVwc(TT0)(0.59)(2260000)Vs=(1000)(250)(4200)(10040)Vs=47250 cm3>Vs0m_sR=m_wc(T-T_0)\\\rho_sV_sR=\rho_wV_wc(T-T_0) \\(0.59)(2260000)V_s=(1000)(250)(4200)(100-40)\\V_s=47250\ cm^3>V_{s0}

Thus, all steam is condensed into water:


Vs=Vs0=150 cm3V_s=V_{s0}=150\ cm^3